# Semantic Operators: A Declarative Model for Rich, AI-based Analytics Over Text Data

### v2

@@ 7. Conclusion
In this work, we proposed semantic operators to provide the first declarative and general-purpose interface to serve bulk-semantic processing. We implement these operators in the LOTUS system to seamlessly extend the relational model and allow programmers to easily compose powerful reasoning-based query pipelines over vast corpora of structured and unstructured data. Our results across a diverse set of applications, including fact-checking, extreme multi-label classification, and search, demonstrate the generality and effectiveness of the semantic operator model as well as new opportunities for efficient algorithms and optimizations. For each task, we find that LOTUS programs capture high quality and state-of-the-art query pipelines with low development overhead, and that they can be automatically optimized with accuracy guarantees to achieve higher performance than existing implementations.
@@ Acknowledgements.
- cus ([n. d.]) [n. d.]. Custom Search JSON API | Programmable Search Engine. https://developers.google.com/custom-search/v1/overview
@@
- gai ([n. d.]) [n. d.]. GAIR-NLP/factool: FacTool: Factuality Detection in Generative AI. https://github.com/GAIR-NLP/factool
- int ([n. d.]) [n. d.]. Introducing Meta Llama 3: The most capable openly available LLM to date. https://ai.meta.com/blog/meta-llama-3/
@@
- ope ([n. d.]) [n. d.]. OpenAI Platform. https://platform.openai.com
@@
- pap ([n. d.]) [n. d.]. Papers with Code - Machine Learning Datasets. https://paperswithcode.com/datasets
@@
- Anantha et al. (2021) Raviteja Anantha, Svitlana Vakulenko, Zhucheng Tu, Shayne Longpre, Stephen Pulman, and Srinivas Chappidi. 2021. Open-Domain Question Answering Goes Conversational via Question Rewriting. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Kristina Toutanova, Anna Rumshisky, Luke Zettlemoyer, Dilek Hakkani-Tur, Iz Beltagy, Steven Bethard, Ryan Cotterell, Tanmoy Chakraborty, and Yichao Zhou (Eds.). Association for Computational Linguistics, Online, 520–534. https://doi.org/10.18653/v1/2021.naacl-main.44
@@
- Gao et al. (2023) Luyu Gao, Zhuyun Dai, Panupong Pasupat, Anthony Chen, Arun Tejasvi Chaganty, Yicheng Fan, Vincent Y. Zhao, Ni Lao, Hongrae Lee, Da-Cheng Juan, and Kelvin Guu. 2023. RARR: Researching and Revising What Language Models Say, Using Language Models. https://doi.org/10.48550/arXiv.2210.08726 arXiv:2210.08726 [cs].
@@
- Izacard et al. (2022) Gautier Izacard, Patrick Lewis, Maria Lomeli, Lucas Hosseini, Fabio Petroni, Timo Schick, Jane Dwivedi-Yu, Armand Joulin, Sebastian Riedel, and Edouard Grave. 2022. Atlas: Few-shot Learning with Retrieval Augmented Language Models. https://doi.org/10.48550/arXiv.2208.03299 arXiv:2208.03299 [cs].
@@
- Khattab et al. (2021) Omar Khattab, Christopher Potts, and Matei Zaharia. 2021. Relevance-guided Supervision for OpenQA with ColBERT. Transactions of the Association for Computational Linguistics 9 (2021), 929–944. https://doi.org/10.1162/tacl_a_00405 Place: Cambridge, MA Publisher: MIT Press.
@@
- Lee et al. (2019) Kenton Lee, Ming-Wei Chang, and Kristina Toutanova. 2019. Latent Retrieval for Weakly Supervised Open Domain Question Answering. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, Anna Korhonen, David Traum, and Lluís Màrquez (Eds.). Association for Computational Linguistics, Florence, Italy, 6086–6096. https://doi.org/10.18653/v1/P19-1612
@@
- Patel et al. (2024) Liana Patel, Peter Kraft, Carlos Guestrin, and Matei Zaharia. 2024. ACORN: Performant and Predicate-Agnostic Search Over Vector Embeddings and Structured Data. Proceedings of the ACM on Management of Data 2, 3 (May 2024), 120:1–120:27. https://doi.org/10.1145/3654923
@@
- Yuksekgonul et al. (2024) Mert Yuksekgonul, Federico Bianchi, Joseph Boen, Sheng Liu, Zhi Huang, Carlos Guestrin, and James Zou. 2024. TextGrad: Automatic ”Differentiation” via Text. http://arxiv.org/abs/2406.07496 arXiv:2406.07496 [cs].
- Zelikman et al. (2022) Eric Zelikman, Yuhuai Wu, Jesse Mu, and Noah D. Goodman. 2022. STaR: Bootstrapping Reasoning With Reasoning. https://doi.org/10.48550/arXiv.2203.14465 arXiv:2203.14465 [cs].
@@
- Zhang et al. (2022) Zhuosheng Zhang, Aston Zhang, Mu Li, and Alex Smola. 2022. Automatic Chain of Thought Prompting in Large Language Models. https://doi.org/10.48550/arXiv.2210.03493 arXiv:2210.03493 [cs].

### v3

@@ 7. Conclusion
In this work, we proposed semantic operators for bulk semantic processing and provide the first formalized model for general-purpose, language-based transformations over data. Our results across a diverse set of applications, including fact-checking, biomedical multi-label classification, search, and topic analysis, demonstrate the generality, expressiveness and robustness of the semantic operator model as well as our optimization approach. For each task, we find that LOTUS programs capture state-of-the-art AI applications with low development overhead, match or exceed result quality of recent LLM-powered analytics systems, and substantially reduce cost, while ensuring accuracy guarantees. Since open-sourcing LOTUS, we have seen a growing user-base, as well as adoption of semantic operators among large database vendors. We believe both represent exciting steps towards new data systems that integrate powerful AI reasoning capabilities over vast knowledge corpora.
@@ Acknowledgements.
- noa ([n.d.]a) [n.d.]a. Custom Search JSON API | Programmable Search Engine. https://developers.google.com/custom-search/v1/overview
@@
- noa ([n.d.]b) [n.d.]b. GAIR-NLP/factool: FacTool: Factuality Detection in Generative AI. https://github.com/GAIR-NLP/factool
- noa ([n.d.]c) [n.d.]c. Introducing Meta Llama 3: The most capable openly available LLM to date. https://ai.meta.com/blog/meta-llama-3/
@@
- noa ([n.d.]d) [n.d.]d. Large Language Model (LLM) Functions (Snowflake Cortex) | Snowflake Documentation. https://docs.snowflake.com/user-guide/snowflake-cortex/llm-functions
@@
- noa ([n.d.]e) [n.d.]e. OpenAI Platform. https://platform.openai.com
@@
- noa ([n.d.]f) [n.d.]f. python-bigquery-dataframes/notebooks/experimental/semantic_operators.ipynb at main · googleapis/python-bigquery-dataframes. https://github.com/googleapis/python-bigquery-dataframes/blob/main/notebooks/experimental/semantic_operators.ipynb
@@
- Anderson et al. (2024) Eric Anderson, Jonathan Fritz, Austin Lee, Bohou Li, Mark Lindblad, Henry Lindeman, Alex Meyer, Parth Parmar, Tanvi Ranade, Mehul A. Shah, Benjamin Sowell, Dan Tecuci, Vinayak Thapliyal, and Matt Welsh. 2024. The Design of an LLM-powered Unstructured Analytics System. https://doi.org/10.48550/arXiv.2409.00847 arXiv:2409.00847 [cs].
@@
- Biswal et al. (2024) Asim Biswal, Liana Patel, Siddarth Jha, Amog Kamsetty, Shu Liu, Joseph E. Gonzalez, Carlos Guestrin, and Matei Zaharia. 2024. Text2SQL is Not Enough: Unifying AI and Databases with TAG. https://doi.org/10.48550/arXiv.2408.14717 arXiv:2408.14717 [cs].
@@
- Codd (1970) E F Codd. 1970. A Relational Model of Data for Large Shared Data Banks. 13, 6 (1970).
- Dai et al. (2024) Hanjun Dai, Bethany Yixin Wang, Xingchen Wan, Bo Dai, Sherry Yang, Azade Nova, Pengcheng Yin, Phitchaya Mangpo Phothilimthana, Charles Sutton, and Dale Schuurmans. 2024. UQE: A Query Engine for Unstructured Databases. https://doi.org/10.48550/arXiv.2407.09522 arXiv:2407.09522 [cs].
@@
- Douze et al. (2024) Matthijs Douze, Alexandr Guzhva, Chengqi Deng, Jeff Johnson, Gergely Szilvasy, Pierre-Emmanuel Mazaré, Maria Lomeli, Lucas Hosseini, and Hervé Jégou. 2024. The Faiss library. https://doi.org/10.48550/arXiv.2401.08281 arXiv:2401.08281 [cs].
@@
- Lewis et al. (2021) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, Sebastian Riedel, and Douwe Kiela. 2021. Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks. https://doi.org/10.48550/arXiv.2005.11401 arXiv:2005.11401 [cs].
@@
- MotherDuck ([n.d.]) MotherDuck. [n.d.]. Introducing the prompt() Function: Use the Power of LLMs with SQL! - MotherDuck Blog. https://motherduck.com/blog/sql-llm-prompt-function-gpt-models/
- Patel et al. (2024a) Liana Patel, Siddharth Jha, Carlos Guestrin, and Matei Zaharia. 2024a. LOTUS: Enabling Semantic Queries with LLMs Over Tables of Unstructured and Structured Data. https://doi.org/10.48550/arXiv.2407.11418 arXiv:2407.11418 [cs] version: 1.
- Patel et al. (2024b) Liana Patel, Peter Kraft, Carlos Guestrin, and Matei Zaharia. 2024b. ACORN: Performant and Predicate-Agnostic Search Over Vector Embeddings and Structured Data. Proceedings of the ACM on Management of Data 2, 3 (May 2024), 120:1–120:27. https://doi.org/10.1145/3654923
@@
- Shankar et al. (2024) Shreya Shankar, Tristan Chambers, Tarak Shah, Aditya G. Parameswaran, and Eugene Wu. 2024. DocETL: Agentic Query Rewriting and Evaluation for Complex Document Processing. https://doi.org/10.48550/arXiv.2410.12189 arXiv:2410.12189 [cs].
@@
- WilliamDAssafMSFT (2024) WilliamDAssafMSFT. 2024. Intelligent Applications - Azure SQL Database. https://learn.microsoft.com/en-us/azure/azure-sql/database/ai-artificial-intelligence-intelligent-applications?view=azuresql
@@
- Zellers et al. (2019) Rowan Zellers, Ari Holtzman, Yonatan Bisk, Ali Farhadi, and Yejin Choi. 2019. HellaSwag: Can a Machine Really Finish Your Sentence? https://doi.org/10.48550/arXiv.1905.07830 arXiv:1905.07830 [cs].
@@
- Zhuang et al. (2024) Honglei Zhuang, Zhen Qin, Kai Hui, Junru Wu, Le Yan, Xuanhui Wang, and Michael Bendersky. 2024. Beyond Yes and No: Improving Zero-Shot LLM Rankers via Scoring Fine-Grained Relevance Labels. http://arxiv.org/abs/2310.14122 arXiv:2310.14122 [cs].